# Statistical Complexity of Heterogeneous Geometric Networks

### v2

@@ V Discussion
Our modelling demonstrated greater statistical complexity arising through the combination of hierarchical and geometric components. While the space of random geometric graphs is regular, the nodes being placed randomly in that space opens up pockets of higher and lower connectivity, as quantified through degree correlations [27]. This may be understood as a classic instance where randomness and regularity interact to generate some degree of complexity. Similarly, heterogeneous random graphs contain an ordered structure in terms of the hierarchy of node degrees, combined with randomness of connections established through the configuration model procedure. By giving a randomly allocated log-normal node fitness to the nodes randomly placed in Euclidean space we see an amplification in terms of complexity. This is particularly interesting since heterogeneous geometric networks closely model many aspects of real-world networks [20]. Future work will explore whether access to heterogeneous connectivity patterns in networks facilitates the formation of more heterogeneous functionality and therefore assess the advantages conferred by the combination of hierarchy and geometry which real-world networks appear to incorporate almost universally.
@@
In the results of HC among the 20 large real-world networks some patterns appear. The general trends among the three groups of networks we have (6 twitch networks, 4 protein networks, 5 physics collaboration networks) indicate that the twitch networks tend to have highest complexity, while protein networks have fairly high complexity and physics collaboration networks have low complexity. The power grid having the lowest complexity can be expected as it is a network with high geometrical constraints and we might expect some specific universal design principles in its construction. On the other hand, online social networks of twitch are largely free from geometrical constraints and may reflect the diversity of social relationships. However, sample sizes would need to be increased to provide stronger evidence for any such generalisations.
The results from the attachment mechanism experiments showed that explanations for the positive relationship between density and hierarchical complexity were not the result of preferential attachment mechanisms for hierarchy or similarity of nodes, nor the combination of the two. This shows us that the increase in HC with respect to density is not a trivial observation. Going further, it points towards a preference of real-world networks towards complexity in novel attachments.
@@
The evidence that random geometric graphs have non-zero HC in the thermodynamic limit gives an interesting insight into the geometrical nature of HC. It tentatively points towards a definitive notion of statistical complexity of networks. Essentially, if we take seriously the established notion of networks as embedded in a latent geometrical space [41, 19], then we can start to conceptualise measures of statistical complexity of networks such as HC as attempting to measure the irregularity of the distribution of points over that space.
@@
In the future, we intend to thoroughly explore the application of normalised HC for analysis of brain networks, across scales, across different types of networks, and eventually, across species. We will also explore applications to protein-protein interaction networks utilising the vast dataset available through SNAP [31] in order to begin to answer questions regarding the relationship between HC of protein-protein interaction networks and evolutionary parameters. Applications to social networks hold obvious appeal given the consistently high complexity we noticed in the twitch social networks. Generally speaking, our understanding of statistical complexity of networks in different fields is limited by a lack of parsimonious measurements for its quantification. The tools offered here can therefore help researchers to begin to answer fundamental questions regarding the existence and extent of statistical complexity of real-world networks, especially in light of the larger and higher quality datasets becoming evermore available.
@@ VI Conclusion
We proposed and demonstrated the utility of a normalisation for hierarchical complexity. We proved that this measure is bounded above by $2$ and tends to zero for Erdös-Rényi random graphs with increasing size. This is analogous to a defining characteristic required of a statistical complexity measure in dynamical systems. We then demonstrated that, while random graph models containing hierarchy and geometry individually had lower complexity, the combined components of hierarchy and geometry is enough to create HC of a similar level to real-world networks. However, real-world networks displayed an association between HC and density that could not be explained solely by these components. Likewise, modelling attachment mechanisms following hierarchy, geometry, or their combination could not explain this association. We conjecture that real-world networks therefore have in-part a preference for attachments which increase complexity of the interacting system. We provide a parsimonious measure for statistical complexity of networks which is ready to be applied to answering questions and gathering new insights into the degrees of complexity in various fields such as neuroscience, protein biology and social networks.
@@ Acknowledgment
The authors would like to thank Dr April S. Kleppe for her thoughtful feedback on our manuscript.

### v3

@@ V Discussion
Our modelling demonstrated greater statistical complexity arising through the combination of hierarchical and geometric components. While the space of random geometric graphs is regular, the nodes being placed randomly in that space opens up pockets of higher and lower connectivity, as quantified through degree correlations [27]. This may be understood as a classic instance where randomness and regularity interact to generate some degree of complexity. Similarly, heterogeneous random graphs contain an ordered structure in terms of the hierarchy of node degrees, combined with randomness of connections established through the configuration model procedure. By giving a randomly allocated log-normal node fitness to the nodes randomly placed in Euclidean space we see an amplification in terms of complexity. This is particularly interesting since heterogeneous geometric networks closely model many aspects of real-world networks [20]. Future work will explore whether access to heterogeneous connectivity patterns in networks facilitates the formation of more heterogeneous functionality and therefore assess the advantages conferred by the combination of hierarchy and geometry which real-world networks appear to incorporate almost universally, as found in e.g. [11] (although there are a subset of networks which buck this trend [32]).
@@
In the results of NHC among the 20 large real-world networks some patterns appear. The general trends among the three groups of networks we have (6 twitch networks, 4 protein networks, 5 physics collaboration networks) indicate that the twitch networks tend to have highest complexity, while protein networks have fairly high complexity and physics collaboration networks have low complexity. The power grid having the lowest complexity can be expected as it is a network with high geometrical constraints and we might expect some specific universal design principles in its construction. On the other hand, online social networks of twitch are largely free from geometrical constraints (although will still have a latent similarity space, but this can be of arbitrarily large dimension) and may reflect the diversity of social relationships. However, sample sizes would need to be increased to provide stronger evidence for any such generalisations.
The results from the link growth mechanism experiments showed that explanations for the positive relationship between density and NHC were not given by link growth mechanisms for popularity or similarity of nodes separately, but again the combination of the two. It indicates that growth of statistical complexity requires a trade-off of randomness and determinism. Too much randomness and no structure is developed. Too little randomness and the structure becomes too rigid. It also highlights how real-world networks may naturally grow to develop higher statistical complexity simply through nodes more likely to (but not with certainty)) form links with nodes with a lot of common neighbours. It also highlights some amount of futility in trying to ever perfectly predict links– there is randomness in connectivity, and in fact networks may well benefit from that randomness in breaking rigidity of patterns and becoming more diverse.
@@
The evidence that random geometric graphs have non-zero NHC in the thermodynamic limit gives an interesting insight into the geometrical nature of NHC. It tentatively points towards a definitive notion of statistical complexity of networks. Essentially, if we take seriously the established notion of networks as embedded in a latent geometrical space [44, 19], then we can start to conceptualise measures of statistical complexity of networks such as NHC as attempting to measure the irregularity of the distribution of points over that space.
It is worth noting that in all of the NHC values we computed, we rarely find anything above 0.1. This means, we expect 0.1 is a very high value of NHC for a network. Here, we recall that a normalisation does not require values to be in any particular range, just that the values are comparable for the same phenomena, for example model with all other parameters fixed, with changes to the targeted normalisation parameter. Further, while we provide an upper bound of 2 to demonstrate there is no possible case of the measure exploding to infinity, we do not believe this is a tight upper bound. From our experience, it is likely a tight upper bound is even below 1.
@@
In the future, we intend to thoroughly explore the application of NHC for analysis of brain networks, across scales, across different types of networks, and eventually, across species. We will also explore applications to protein-protein interaction networks utilising the vast datasets available through the STRING database in order to begin to answer questions regarding the relationship between NHC of protein-protein interactions and evolutionary parameters. Applications to social networks hold obvious appeal given the consistently high complexity we noticed in the twitch social networks. There is also serious scope for extension and improvement of NHC. For example, we could consider generalising the measure to considering neighbourhoods of neighbourhoods, up to an arbitrary depth. Generally speaking, our understanding of statistical complexity of networks in different fields is limited by a lack of parsimonious measurements for its quantification. The tools offered here can therefore help researchers to begin to answer fundamental questions regarding the existence and extent of statistical complexity of real-world networks, especially in light of the larger and higher quality datasets becoming evermore available.
@@ VI Conclusion
We proposed and demonstrated the utility of a normalisation for hierarchical complexity– NHC. We proved that this measure is bounded above by $2$ and tends to zero for Erdös-Rényi random graphs with increasing size. This is analogous to a defining characteristic required of a statistical complexity measure in dynamical systems. We then demonstrated that, while random graph models containing degree heterogeneity and geometry individually had lower complexity, the combined components of degree heterogeneity and geometry is enough to create NHC of a similar level to real-world networks. However, we then found that real-world networks displayed an association between NHC and density that could not be explained solely by our models. Instead, we could manage to explain this consistently with a common neighbours link growth algorithm with exponentiated probabilities. Particularly, this was more consistent than a deterministic weight ranking algorithm and the more random non-exponentiated probability algorithm. All other algorithms tried failed to increase complexity. We therefore posit that real-world networks have a preference for growth which increases complexity of the interacting system. We provide a parsimonious measure for statistical complexity of networks which is ready to be applied to answering questions and gathering new insights into the degrees of complexity in various fields such as neuroscience, protein biology and social networks.
@@ Acknowledgment
The authors would like to thank Dr April S. Kleppe for her thoughtful feedback on our work.